OpenStudy (anonymous):

In H5 the parser don't "like" my formulas. What's wrong with them: vin*(1-1/sqrt(1+2K*RS*(VIN-VT)))) and: (vIN-VT)/RS + (1/K*RS^2)*(1-sqrt(1+2K*RS*(vIN-VT)))

OpenStudy (anonymous):

Worked, only one left :)

OpenStudy (anonymous):

This one left - (vIN-VT)/RS + (1/K*RS^2)*(1-sqrt(1+2K*RS*(vIN-VT))), now there isn't a problem with parsing, but seem the formula isn't correct.

OpenStudy (anonymous):

\[(vin*(1-1/\sqrt(1+2*K*RS*(VIN-VT))))\] does not work and i cannot see why

OpenStudy (anonymous):

that is the th equation i wrote

OpenStudy (anonymous):

These formulas are for H5P3 and it's correct. I need another one: Q3 for H5P2 SOURCE FOLLOWER LARGE SIGNAL

OpenStudy (anonymous):

I got it!!! Actually in finding valid Operating ranges \[V0 = [-1 + \sqrt{1+2*K*RL*Vs} ] \div KRL\] In the above formula Vs= applied voltage across RL and Drain to Source( i.e Vds).. \[Vs=ids∗RL+Vds(or)Vs=ids∗RL+V0\] But in given problem , by applying KVL to output loop we get \[Vds+idsRL=Vs+(−Vs−)\] \[Vds+ids∗RL=1−(−1)=2\] So in finding V0 substitute\[Vs=2 \] (NOT 1) And we know Valid maximum input range is \[Vgs = V0 + Vthershold\] Applying KVL to input loop \[Vin=Vgs+Vs−\] Substitue the value ...You will get it !!!!!! Cheers!!!!

OpenStudy (benzo):

help please h5p2 question# 7....What is the minimum value of VDD (in Volts) that we must supply to keep the transistor in the region of saturated operation?

OpenStudy (anonymous):

Conditions for Saturation region Opereation are.... \[VGS \ge VT\]\[VDS \ge VGS-VT\]For the given Circuit \[VDS = VDD-VOUT\]\[VGS=VIN-VOUT\] By using all conditions we get final Condition ....\[VDD \ge VIN-VT\]Here Substitue Maximum Possible \[VIN\] (As given in your problem.i.e maximum positive swing) Now you have \[VIN\] and \[VT\] So subtitue in the Final Condition for \[VDD \] ,You will get the Solution... Cheers!!!!

OpenStudy (benzo):

help please h5p2 question# 7....What is the minimum value of VDD (in Volts) that we must supply to keep the transistor in the region of saturated operation?

OpenStudy (anonymous):

@benzo I gave an clear explanation above to find VDD ....are you interested in knowing answer or the process of getting answer ?... Anyway for the numbers given in my problem set Max \[VIN=7.1\] and Minimun \[VIN=5.7\] So I got Minimum Possible \[VDD=5.1\]

OpenStudy (benzo):

Srikanth_Gangula i got it thanks lab and home work 100%

OpenStudy (anonymous):

You are Welcome Benzo!!!!